# AdapDISCOM: An Adaptive Sparse Regression Method for High-Dimensional Multimodal Data With Block-Wise Missingness and Measurement Errors

### v1

@@ 5 Application to Multimodal High-Dimensional Data from the ADNI Cohort
Alzheimer’s disease represents a major global public health challenge. In 2018, Alzheimer’s Disease International estimated that approximately 50 million individuals worldwide were living with dementia [20], a number projected to triple by 2050 [38]. Advancing research on Alzheimer’s disease is thus critical to mitigating its impact on population health and quality of life. In this context, the Alzheimer’s Disease Neuroimaging Initiative (ADNI) [36] has played a pivotal role.
@@
We applied the AdapDISCOM family of methods to evaluate the association between multimodal ADNI data and Mini-Mental State Examination (MMSE) scores, while accounting for the complex structure of missingness and measurement error. MMSE is a widely used 30-point questionnaire designed to assess cognitive function, particularly in older adults or patients suspected of dementia [17]. Specifically, we analyzed data from four modalities: MRI-derived volumetric measures, Flortaucipir PET SUVRs, CSF biomarkers, and genetic variants (SNPs).
@@
  • Scenario I: three modalities (CSF + MRI + PET), analogous to the setup in Yu et al., [54], excluding genetic data.
@@
In Scenario I, 196 subjects had complete data (split into 116 for testing, 40 for tuning, and the remaining 40 plus 1,015 incomplete cases for training), with 53.6% missingness overall. In Scenario II, only 86 subjects had complete data (split into 40 for testing, 36 for tuning, and 10 plus 1,125 incomplete cases for training), with 53.44% missingness. For each scenario, hyperparameters were tuned using a 30-point grid based on KKT conditions to define $\lambda_{\max},$ with $\lambda_{\min}$ set as a fraction $(10^{-2})$ of $\lambda_{\max}$ following the GLMnet approach [18]. Each method was run 40 times per scenario, with performance assessed via MSE and $R^{2}$ on the test set.
@@ 5.1 Results
[…]
@@ 6 Discussion
Several avenues remain open for further extending and generalizing AdapDISCOM. One natural direction is to adapt the method to generalized linear models and longitudinal data, and to explore alternative penalty functions beyond the standard LASSO. Since AdapDISCOM was originally developed under the missing completely at random (MCAR) assumption, another critical area of future research is its extension to settings with missing at random (MAR) or missing not at random (MNAR) mechanisms. Beyond measurement error and missing data, AdapDISCOM could also be enhanced to detect outliers, inspired by the approach in [2]. Finally, incorporating expectiles and M-quantiles [3] offers a promising direction, enabling inference beyond the mean and capturing more nuanced features of the response distribution.
@@ 10.1 Proof
In this section, we present the proofs of the results stated in the manuscript, which are a generalization of the DISCOM method. With close attention, it can be seen that, with very few exceptions, the proofs of the generalization are identical to those presented by Yu et al., [54]. We add them here for completeness.
@@
By minimizing the expression $\sum_{k=1}^{K}(1-\alpha_{k})^{2}\|(\gamma\mathbb{I}_{p}-\Sigma_{I_{k}})\|^{2}_{F}$ with respect to $gamma$ we get the optimal value: $\gamma*=\frac{1}{p}\sum_{k=1}^{K}\frac{(1-\alpha_{k}^{*})^{2}}{\sum_{t=1}^{K}(1-\alpha_{t}^{*})^{2}}\operatorname{Tr}(\widetilde{\Sigma}_{I_{k}}).$ The optimal value of $\alpha_{C}$ is obtained by minimizing $\operatorname{\mathbb{E}}[\|(\alpha_{C}\widetilde{\Sigma}_{C}-\Sigma_{C})\|^{2}_{F}]$ and then we have $\alpha_{C}^{*}=\frac{\|\Sigma_{C}\|^{2}_{F}}{\|\Sigma_{C}\|^{2}_{F}+\delta_{C}^{2}}.$ Replacing $\gamma$ by its optimal value in the objective function and taking the derivative with respect to $\alpha_{k},$ we obtain, for $k\in\{1,2,\ldots,K\},$ the optimal value $\alpha_{k}^{*}=\frac{\theta_{I_{k}}^{2}}{\theta_{I_{k}}^{2}+\delta_{I_{k}}^{2}}.$ Thus, the optimal value of $\alpha_{p}$ is $\alpha_{p}^{*}=\gamma^{*}\sum_{k=1}^{K}(1-\alpha_{k}^{*})=\gamma^{*}\sum_{k=1}^{K}\frac{\delta_{I_{k}}^{2}}{\theta_{I_{k}}^{2}+\delta_{I_{k}}^{2}}.$
@@
Ravikumar et al., [40] Consider a zero-mean random vector $(X_{1},\ldots,X_{p})$ with covariance $\Sigma$ such that each $X_{j}/\sqrt{\sigma_{jj}}$ is sub-Gaussian with parameter L. Given n i.i.d. samples, the associated sample covariance $\widehat{\Sigma}$ satisfies the following tail bound
@@
$\widetilde{\text{C}}.$   The random variable $y/\sqrt{\operatorname{\mathbb{V}\text{ar}}(y)}$ is sub-Gaussian with parameter $\frac{L}{\min\{1,\sqrt{\operatorname{\mathbb{V}\text{ar}}(y)}\}}$ and accordingly the upper bound of $\delta$ in Lemma 1 is $8\max\{1,\operatorname{\mathbb{V}\text{ar}}(y)\}\frac{L^{2}}{\min\{1,\sqrt{\operatorname{\mathbb{V}\text{ar}}(y)}\}}.$ Now taking $\delta=\nu_{3}\sqrt{\frac{\log p}{n_{j}}},\,\nu_{3}=16(1+4\frac{L^{2}}{\min\{1,\operatorname{\mathbb{V}\text{ar}}(y)\}})\max\{1,\operatorname{\mathbb{V}\text{ar}}(y)\},\;\nu_{4}=4$ and obsserving that $\min_{j}n_{j}\geq min_{j,t}n_{jt}>6logp>4logp$, then applying Lemma 1 gives as:
@@
Now consider, $\widehat{\Sigma}=\sum_{k=1}^{K}\alpha_{k}\widetilde{\Sigma}_{I_{k}}+\alpha_{C}\widetilde{\Sigma}_{C}+\sum_{k=1}^{K}(1-\alpha_{k})\mathbb{I}_{p},$ where $1-\alpha_{k}=O(\sqrt{\log p/\min_{j}n_{j}}),\;k\in\{1,2,\ldots K\}$ and $1-\alpha_{C}=O(\sqrt{\log p/\min_{j,t}n_{jt}}).$ Denote events $\mathcal{A}=\{\|\widehat{\Sigma}-\Sigma\|_{\max}\geq\nu_{1}\prime\sqrt{\log p/\min_{j,t}n_{jt}}\}$ and =$\mathcal{B}=\{\|\widetilde{\text{C}}-\text{C}\|_{\max}\geq\nu_{3}\sqrt{\log p/\min_{j}n_{j}}\}.$ Based on the definition of $\widehat{\Sigma}=(\sigma_{jt})_{j,t=1}^{p},$ we know that
$\widehat{\sigma}_{jt}-\sigma_{jt}=\begin{cases}\sum_{k\neq k^{\prime}}^{K}(1-\alpha_{k})&\text{if $j=t$ and $j\in$ modality $k^{\prime}$;}\\ \alpha_{k}\widetilde{\sigma}_{jt}-\sigma_{jt}&\text{if $j\neq t$ \ ($j$ and $t$ are in the same modality);}\\ \alpha_{C}\widetilde{\sigma}_{jt}-\sigma_{jt}&\text{if $j\neq t$ \ ($j$ and $t$ are in different modalities).}\\ \end{cases}$
@@
$\begin{split}\lvert\widehat{\sigma}_{jt}-\sigma_{jt}\rvert=\lvert\alpha_{k}\widetilde{\sigma}_{jt}-\sigma_{jt}\rvert&\leq\alpha_{k}\lvert\widetilde{\sigma}_{jt}-\sigma_{jt}\rvert+(1-\alpha_{k})\lvert\sigma_{jt}\rvert\\ &\leq\alpha_{k}\lvert\widetilde{\sigma}_{jt}-\sigma_{jt}\rvert+(1-\alpha_{k})\\ &\leq\alpha_{k}\nu_{1}\sqrt{\log p/\min_{j}n_{j}}+(1-\alpha_{k})\quad\text{derived from above results}\\ &\leq\nu_{1}\sqrt{\log p/\min_{j}n_{j}}+(1-\alpha_{k}).\\ \end{split}$
@@
[…]
@@
for sufficiently large $s,\ p,$ and $\min_{j,t}n_{jt}.$ On the other hand, we have
@@
$\breve{\text{C}}.$ Noting that: for each $j\in\{1,2,\ldots,p\}$ and $i\in S_{j},$ we have $\operatorname{\mathbb{V}\text{ar}}(x_{ij}y_{i})\leq\operatorname{\mathbb{E}}(x_{ij}^{2}y_{i}^{2})=\sqrt{\operatorname{\mathbb{E}}(x_{ij}^{4})\operatorname{\mathbb{E}}(y_{i}^{4})}\leq 2(Q_{1}+Q_{2})^{2}.$ Now noting, $\text{H}_{j}=(\text{Q}_{1}+\text{Q}_{2})\sqrt{n_{j}/\log p}$ for each $j\in\{1,2,\ldots,p,$ and using Theorem 5 in Fan et al., [13], we have
@@ 10.2.1 Preprocessing and Quality Control of MRI Data
  • Brain extraction was conducted using the ANTs antsBrainExtraction.sh method [1, 48].
  • Tissue segmentation of gray matter, white matter, and cerebrospinal fluid was performed using FSL FAST [56].
  • Spatial normalization of gray matter, white matter, and cerebrospinal fluid was performed using FSL FAST to the MNI152 template was carried out with ANTs registration algorithms [1, 48].
@@
  • Bias field correction, using the N4ITK algorithm [47].
  • Segmentation of subcortical white matter and deep gray matter structures, according to the Aseg atlas [15, 16].
  • Tessellation of the gray/white matter boundary and cortical surface reconstruction [7, 14].
@@
  • Regional cortical thickness (CTh) measurements in each region of the Desikan-Killiany (DK) atlas [9].
  • Regional subcortical volumes, as defined by the Aseg atlas [15].
@@ 10.2.2 Preprocessing and Quality Control of Flortaucipir PET (SUVRs)
Data acquisition and initial preprocessing. We downloaded raw flortaucipir PET images in their fully preprocessed format from the ADNI repository (series description: AV1451 Coreg, Avg, Std Img and Vox Siz, Uniform Resolution). These images had already undergone standard ADNI preprocessing, including spatial standardization and initial intensity normalization to a cerebellar cortex region [25]. However, following UC Berkeley recommendations, we replaced this initial normalization with a subject-specific normalization based on FreeSurfer-defined reference regions in native space to reduce warping-induced noise and improve regional specificity.
@@
SUVR computation. SUVRs (standardized uptake value ratios) were computed in native space within FreeSurfer-defined ROIs. Specifically, regional flortaucipir uptake values were normalized to the inferior cerebellar grey matter, identified using a combination of SUIT cerebellar atlas regions and FreeSurfer cerebellar grey masks, excluding regions prone to off-target binding [10]. This reference region is preferred to mitigate dorsal cerebellar contamination and improve comparability across subjects.
For regional quantification, we used ROIs defined by FreeSurfer segmentation, including cortical and subcortical areas, as well as composite regions approximating Braak stages (Braak 1–6) and a meta-temporal composite ROI [31, 42]. SUVRs were calculated as the mean uptake within each target ROI divided by the mean uptake in the inferior cerebellar grey reference region.
Partial volume correction. To mitigate partial volume effects, we also computed SUVRs corrected using the Geometric Transfer Matrix (GTM) method [41], modelling both ROIs and off-target binding regions such as the choroid plexus [31]. Both PVC and non-PVC SUVR datasets were generated, with the non-PVC dataset using MRI scans closest in time to each PET scan and the PVC dataset using a baseline MRI across all time points.

### v2

@@ 5 Application to Multimodal High-Dimensional Data from the ADNI Cohort
Alzheimer’s disease represents a major global public health challenge. In 2018, Alzheimer’s Disease International estimated that approximately 50 million individuals worldwide were living with dementia [Gustavsson et al., 2023], a number projected to triple by 2050 [Prince et al., 2015]. Advancing research on Alzheimer’s disease is thus critical to mitigating its impact on population health and quality of life. In this context, the Alzheimer’s Disease Neuroimaging Initiative (ADNI) [Mueller et al., 2005a] has played a pivotal role.
@@
We applied the AdapDISCOM family of methods to evaluate the association between multimodal ADNI data and Mini-Mental State Examination (MMSE) scores, while accounting for the complex structure of missingness and measurement error. MMSE is a widely used 30-point questionnaire designed to assess cognitive function, particularly in older adults or patients suspected of dementia [Folstein et al., 1975]. Specifically, we analyzed data from four modalities: MRI-derived volumetric measures, Flortaucipir PET SUVRs, CSF biomarkers, and genetic variants (SNPs).
@@
  • Scenario I: three modalities (CSF + MRI + PET), analogous to the setup in Yu et al. [2020], excluding genetic data.
@@
In Scenario I, 196 subjects had complete data (split into 116 for testing, 40 for tuning, and the remaining 40 plus 1,015 incomplete cases for training), with 53.6% missingness overall. In Scenario II, only 86 subjects had complete data (split into 40 for testing, 36 for tuning, and 10 plus 1,125 incomplete cases for training), with 53.44% missingness. For each scenario, hyperparameters were tuned using a 30-point grid based on KKT conditions to define $\lambda_{\max},$ with $\lambda_{\min}$ set as a fraction $(10^{-2})$ of $\lambda_{\max}$ following the GLMnet approach [Friedman et al., 2010]. Each method was run 40 times per scenario, with performance assessed via MSE and $R^{2}$ on the test set.
@@ 5.1 Results
Figure 4 illustrates variables selected consistently $(\geq 50\%,\geq 75\%,\ \text{and}\ \geq 100\%$ of the time) by each method, along with their effect sizes (color gradient). Results were consistent across scenarios, with slightly more robustly selected variables in Scenario I. We observed two distinct groups of methods: AdapDISCOM/DISCOM and LASSO-based/SCOM. The variables selected by AdapDISCOM and DISCOM differed from those selected by LASSO-based methods and SCOM, which tended to select more variables with negligible or null effects. In contrast, variables selected consistently by AdapDISCOM and DISCOM had substantial, directionally consistent effects across methods, further demonstrating the inefficacy of conventional approaches in the presence of measurement error and missing data.
@@ 6 Discussion
Several avenues remain open for further extending and generalizing AdapDISCOM. One natural direction is to adapt the method to generalized linear models and longitudinal data, and to explore alternative penalty functions beyond the standard LASSO. Since AdapDISCOM was originally developed under the missing completely at random (MCAR) assumption, another critical area of future research is its extension to settings with missing at random (MAR) or missing not at random (MNAR) mechanisms. Beyond measurement error and missing data, AdapDISCOM could also be enhanced to detect outliers, inspired by the approach in [Barry et al., 2022]. Finally, incorporating expectiles and M-quantiles [Barry et al., 2023] offers a promising direction, enabling inference beyond the mean and capturing more nuanced features of the response distribution.
@@ 10.1 Proof
In this section, we present the proofs of the results stated in the manuscript, which are a generalization of the DISCOM method. With close attention, it can be seen that, with very few exceptions, the proofs of the generalization are identical to those presented by Yu et al. [2020]. We add them here for completeness.
@@
By minimizing the expression $\sum_{k=1}^{K}(1-\alpha_{k})^{2}\|(\gamma\mathbb{I}_{p}-\Sigma_{I_{k}})\|^{2}_{F}$ with respect to gammagamma we get the optimal value: $\gamma*=\frac{1}{p}\sum_{k=1}^{K}\frac{(1-\alpha_{k}^{*})^{2}}{\sum_{t=1}^{K}(1-\alpha_{t}^{*})^{2}}\operatorname{Tr}(\widetilde{\Sigma}_{I_{k}}).$ The optimal value of $\alpha_{C}$ is obtained by minimizing $\operatorname{\mathbb{E}}[\|(\alpha_{C}\widetilde{\Sigma}_{C}-\Sigma_{C})\|^{2}_{F}]$ and then we have $\alpha_{C}^{*}=\frac{\|\Sigma_{C}\|^{2}_{F}}{\|\Sigma_{C}\|^{2}_{F}+\delta_{C}^{2}}.$ Replacing $\gamma$ by its optimal value in the objective function and taking the derivative with respect to $\alpha_{k},$ we obtain, for $k\in\{1,2,\ldots,K\},$ the optimal value $\alpha_{k}^{*}=\frac{\theta_{I_{k}}^{2}}{\theta_{I_{k}}^{2}+\delta_{I_{k}}^{2}}.$ Thus, the optimal value of $\alpha_{p}$ is $\alpha_{p}^{*}=\gamma^{*}\sum_{k=1}^{K}(1-\alpha_{k}^{*})=\gamma^{*}\sum_{k=1}^{K}\frac{\delta_{I_{k}}^{2}}{\theta_{I_{k}}^{2}+\delta_{I_{k}}^{2}}.$
@@
Ravikumar et al. [2011] Consider a zero-mean random vector $(X_{1},\ldots,X_{p})$ with covariance $\Sigma$ such that each $X_{j}/\sqrt{\sigma_{jj}}$ is sub-Gaussian with parameter L. Given n i.i.d. samples, the associated sample covariance $\widehat{\Sigma}$ satisfies the following tail bound
@@
[…]
@@
Now consider, $\widehat{\Sigma}=\sum_{k=1}^{K}\alpha_{k}\widetilde{\Sigma}_{I_{k}}+\alpha_{C}\widetilde{\Sigma}_{C}+\sum_{k=1}^{K}(1-\alpha_{k})\mathbb{I}_{p},$ where $1-\alpha_{k}=O(\sqrt{\log p/\min_{j}n_{j}}),\;k\in\{1,2,\ldots K\}$ and $1-\alpha_{C}=O(\sqrt{\log p/\min_{j,t}n_{jt}}).$ Denote events $\mathcal{A}=\{\|\widehat{\Sigma}-\Sigma\|_{\max}\geq\nu_{1}\prime\sqrt{\log p/\min_{j,t}n_{jt}}\}$ and $\mathcal{B}=\{\|\widetilde{\text{C}}-\text{C}\|_{\max}\geq\nu_{3}\sqrt{\log p/\min_{j}n_{j}}\}.$ Based on the definition of $\widehat{\Sigma}=(\sigma_{jt})_{j,t=1}^{p},$ we know that
[…]
@@
[…]
@@
Therefore, we have $\|\widetilde{\text{C}}-\widehat{\Sigma}\boldsymbol{\beta}^{0}\|_{\max}=O_{p}\Big(\|\boldsymbol{\beta}^{0}\|_{1}\sqrt{\log p/\min_{j,t}n_{jt}}\Big).$ Furthermore, under the condition $(\textbf{A2}),$ if the event $\mathcal{A}$ occurs, we have
@@
for sufficiently large s,p,s,\ p, and $\min_{j,t}n_{jt}.$ On the other hand, we have
@@
[…]
@@ 10.2.1 Preprocessing and Quality Control of MRI Data
  • Brain extraction was conducted using the ANTs antsBrainExtraction.sh method [Avants et al., 2008; Tustison et al., 2021].
  • Tissue segmentation of gray matter, white matter, and cerebrospinal fluid was performed using FSL FAST [Zhang et al., 2002].
  • Spatial normalization of gray matter, white matter, and cerebrospinal fluid was performed using FSL FAST to the MNI152 template was carried out with ANTs registration algorithms [Avants et al., 2008; Tustison et al., 2021].
@@
  • Bias field correction, using the N4ITK algorithm [Tustison et al., 2010].
  • Segmentation of subcortical white matter and deep gray matter structures, according to the Aseg atlas [Fischl et al., 2002, 2004].
  • Tessellation of the gray/white matter boundary and cortical surface reconstruction [Dale et al., 1999; Fischl and Dale, 2000].
@@
  • Regional cortical thickness (CTh) measurements in each region of the Desikan-Killiany (DK) atlas [Desikan et al., 2006].
  • Regional subcortical volumes, as defined by the Aseg atlas [Fischl et al., 2002].
@@ 10.2.2 Preprocessing and Quality Control of Flortaucipir PET (SUVRs)
Data acquisition and initial preprocessing. We downloaded raw flortaucipir PET images in their fully preprocessed format from the ADNI repository (series description: AV1451 Coreg, Avg, Std Img and Vox Siz, Uniform Resolution). These images had already undergone standard ADNI preprocessing, including spatial standardization and initial intensity normalization to a cerebellar cortex region [Jagust et al., 2015]. However, following UC Berkeley recommendations, we replaced this initial normalization with a subject-specific normalization based on FreeSurfer-defined reference regions in native space to reduce warping-induced noise and improve regional specificity.
@@
SUVR computation. SUVRs (standardized uptake value ratios) were computed in native space within FreeSurfer-defined ROIs. Specifically, regional flortaucipir uptake values were normalized to the inferior cerebellar grey matter, identified using a combination of SUIT cerebellar atlas regions and FreeSurfer cerebellar grey masks, excluding regions prone to off-target binding [Diedrichsen, 2006]. This reference region is preferred to mitigate dorsal cerebellar contamination and improve comparability across subjects.
For regional quantification, we used ROIs defined by FreeSurfer segmentation, including cortical and subcortical areas, as well as composite regions approximating Braak stages (Braak 1–6) and a meta-temporal composite ROI [Maass et al., 2017; Schöll et al., 2016]. SUVRs were calculated as the mean uptake within each target ROI divided by the mean uptake in the inferior cerebellar grey reference region.
Partial volume correction. To mitigate partial volume effects, we also computed SUVRs corrected using the Geometric Transfer Matrix (GTM) method [Rousset et al., 1998], modelling both ROIs and off-target binding regions such as the choroid plexus [Maass et al., 2017]. Both PVC and non-PVC SUVR datasets were generated, with the non-PVC dataset using MRI scans closest in time to each PET scan and the PVC dataset using a baseline MRI across all time points.